OpenStudy (anonymous):

Which of the following is the solution to -x − 10 = -3?

OpenStudy (anonymous):

-x - 10 = -3 -x = 7 x = -7

OpenStudy (anonymous):

The solution to 4x + 1 = 15 is x = 4. True of False?

OpenStudy (anonymous):

try it for yourself and ill correct it for you

OpenStudy (anonymous):

and show me your working

OpenStudy (anonymous):

I think it's false

OpenStudy (anonymous):

4*4 =16 +1 equals 17 not 15

OpenStudy (anonymous):

Can you help me with the next question

OpenStudy (anonymous):

go to www.wolframalpha.com

OpenStudy (anonymous):

u will get steps along with answer :)